OpenStudy (anonymous):

simplify the given expression 8+2^3

OpenStudy (anonymous):

8+(2)(2)(2) = 8+(4)(2)=8+8=16

OpenStudy (anonymous):

You solve the exponent first (PEMDAS). So 2^3 is 8. 8+8=16

OpenStudy (anonymous):

use parenthesis... 8+2^3=8+8=16